OpenStudy (anonymous):

I don't really remember how to solve this can someone help walk me through it? 2x^2=-7x+4 I need to solve by factoring..

pooja195 (pooja195):

Ok basically first we need to put the equation into standered form which looks like this: \[\huge~\rm~\bf~ax^2+bx+c=0\] Can you you do that please?

OpenStudy (anonymous):

Okay If I do that I get 2x^2+7x-4=0

pooja195 (pooja195):

Ok good \[\huge~\rm~\bf2x^2+7x−4=0\] Lets use factor by grouping method 

pooja195 (pooja195):

So what is \[\huge~\rm~-4 \times 2=? \]

OpenStudy (anonymous):

-8

pooja195 (pooja195):

Good the 7 remains the same now we have 

pooja195 (pooja195):

what NUMBERS multiply to -8 but add/subtract to make 7?

OpenStudy (anonymous):

-7 & -1

OpenStudy (anonymous):

Nevermind. I did that backwards

pooja195 (pooja195):

Those 2 aren't right -7 x -1=7 -7-1=-8

OpenStudy (anonymous):

8 & 1

pooja195 (pooja195):

Good almost there it would be 8 and -1

OpenStudy (anonymous):

Yup! (:

pooja195 (pooja195):

7x has been split into -1 and 8 now our equation looks like this \[\huge~\rm~\bf2x^2-1+8x-4\] \[\huge~\rm~\bf(2x^2-1)(8x-4)\] We need to find a common factor for both of these

OpenStudy (anonymous):

When I factored it out I got (2x-1)(x+4)

pooja195 (pooja195):

That's the correct answer :)

OpenStudy (anonymous):

So now I just set them each to zero and I should have both zeros right?